OpenStudy (anonymous):

How many liters of oxygen gas, at standard temperature and pressure, will react with 35.4 grams of calcium metal? Show all of the work used to solve this problem. 2 Ca + O2 ---> 2 CaO

OpenStudy (joannablackwelder):

This one is pretty similar.Can you see how far you can get?

OpenStudy (anonymous):

Ok so we are gonna use the same formula right?

OpenStudy (joannablackwelder):

You got it.

OpenStudy (anonymous):

Standard temp = 273K Standard Pressure= 1atm

OpenStudy (anonymous):

PV=nRT P=1 R=0.0821 T=273

OpenStudy (joannablackwelder):

Pretty close to 1 atm. It is technically 1 bar which is 0.987 atm, but close enough.

OpenStudy (anonymous):

Ok and would n= 2?

OpenStudy (joannablackwelder):

I don't get that.

OpenStudy (anonymous):

I dont know what to do from here

OpenStudy (joannablackwelder):

Use the grams of Ca to get moles of Ca.

OpenStudy (joannablackwelder):

35.4 g Ca*(1 mol/40 g Ca)=.885 mol Ca

OpenStudy (anonymous):

ohh ok

OpenStudy (joannablackwelder):

.885 mol Ca*(1 mol O2/2 mol Ca)=

OpenStudy (joannablackwelder):

0.4425 mol O2

OpenStudy (anonymous):

.4425/1.77?

OpenStudy (joannablackwelder):

What does that get?

OpenStudy (anonymous):

.25 Would that be n?

OpenStudy (joannablackwelder):

Why did you divide by 1.77?

OpenStudy (anonymous):

because if one mol of Ca=.885 and you have two that would be 1.77mol right?

OpenStudy (joannablackwelder):

One mole of Ca isn't .885. You have .885 moles of Ca is what we have to start with.

OpenStudy (anonymous):

Yeh im sorry im getting confused

OpenStudy (anonymous):

so would it just be .4425/.885

OpenStudy (joannablackwelder):

Now I'm confused why you would do that...

OpenStudy (anonymous):

ok lets star from = .885 mol Ca*(1 mol O2/2 mol Ca)=

OpenStudy (joannablackwelder):

There you go!!! :D

OpenStudy (joannablackwelder):

How many moles of O2 is that?

OpenStudy (anonymous):

Just one?

OpenStudy (joannablackwelder):

Why?

OpenStudy (anonymous):

I dont know thats as far as i can go

OpenStudy (joannablackwelder):

.885 mol Ca*(1 mol O2/2 mol Ca)=0.4425 mol O2

OpenStudy (joannablackwelder):

You had it, you just didn't simplify.

OpenStudy (anonymous):

Ok but can you explain how you got 0.4425 mol O2

OpenStudy (joannablackwelder):

.885 mol Ca*(1 mol O2/2 mol Ca) means you take 0.885 *1/2=0.4425

OpenStudy (anonymous):

Ohhhhh ok i got it

OpenStudy (joannablackwelder):

mol Ca*mol O2/mol Ca = mol O2 :)

OpenStudy (joannablackwelder):

Yay!

OpenStudy (anonymous):

so now we have n right?

OpenStudy (joannablackwelder):

You got it!

OpenStudy (anonymous):

P=1 R=0.0821 T=273 n=4425 V=x

OpenStudy (joannablackwelder):

n=0.4425

OpenStudy (anonymous):

right sorry

OpenStudy (joannablackwelder):

No worries :)

OpenStudy (anonymous):

V=9.917?

OpenStudy (joannablackwelder):

That's what I get :)

OpenStudy (anonymous):

awsome

OpenStudy (joannablackwelder):

in Liters

OpenStudy (joannablackwelder):

Yup!

OpenStudy (anonymous):

Ok im getting it now just need two more

OpenStudy (anonymous):

No.. This is not a practice exam. This is a module quiz from FLVS so using this answer will be an academic integrity violation. Plus it has a obvious mistake in it. Mr. Carlyle FLVS Chemistry Instructor This question is a violation of the OpenStudy Guidelines. The question is from an online school plus a violation by solicitation for members to unknowingly assist them cheating on an exam. 30653065chem30653065